ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 60, 62, 65 (65C10, 65U05), 68 (68-04) Paditz, Ludwig, HTW Dresden(FH) - Univ. of appl. science, Germany: Comparision of the statistic calculators TI-83, EL-9600 and CFX-9850G PLUS. By the help of a program-example (“growing dandelions′′) several graphic calculators with additional statistic functions are compared. The random number generator is used (random scatter-plot). The simulation will be analysed in a histogram and by the help of the Binomial- and Poissondistribution. Reference: B. Dye, Growing dandelions on an EL9600, Internet:

http://www.anglia.co.uk/education/mathsnet/graphcal/dandelion EL9600. html

L. Paditz, Zufallsaussaat und Wachstum von Loewenzahn (“Pusteblumen”),

Internet: http://www.informatik.htw-dresden.de/˜ paditz/padite96.html